OpenStudy (tia_sweetheart17):

solve the equation: x(7x-2)=9

OpenStudy (rina.r):

x(7x-2)=9 7x2-2x=9 (its 7x square) 5x =9 x= 5/9 x=0.5

OpenStudy (anonymous):

no we can't subtract x from 7xsquare

OpenStudy (rina.r):

no u multiply x with the number inside the bracket which gives u x square

OpenStudy (rina.r):

x come with the number when u subtract 7x from -2x

OpenStudy (anonymous):

i am not sure so use other means for confirm